what is the area bounded by the curves y=8-x^2.


1
Expert's answer
2021-05-24T08:22:02-0400

"y=8-x^2"

"0=8-x^2"

"x=2.82842\,\:x=-2.82842"

"\int^{2.8284}_{-2.8284}(8-x^2)dx"

"30.16988" units sq